OpenStudy (anonymous):

Express the sum using summation notation. Use 1 as the lower limit of summation and i for the index summation. 1/3+1/2+3/5+...+5/6

OpenStudy (freckles):

what do you notice about the denominators 1/3, 2/4,3/5,4/6,5/7,6/8,7/9,8/10,9/11,10/12

OpenStudy (anonymous):

They increase by 1

OpenStudy (freckles):

\[\frac{1}{3},\frac{2}{4},\frac{3}{5},\frac{4}{6},\frac{5}{7},\frac{6}{8},\frac{7}{9},\frac{8}{10},\frac{9}{11},\frac{10}{12}\] yeah in the numerators are also being increased by one

OpenStudy (freckles):

10/12=5/6 by the way

OpenStudy (freckles):

and your 1/2 from before equals 2/4

OpenStudy (anonymous):

Ohhhhh, okay. THANK YOU!

OpenStudy (freckles):

so the sequence I'm looking at relates to your series

OpenStudy (freckles):

do you know how to come up with summation notation thingy now?

OpenStudy (anonymous):

Yes. I was just having a hard time coming up with the equation.

OpenStudy (freckles):

cool stuff! it was kinda hard to see at first like I was hmm how is that 5/6 going to come into play and I notice 1/3 and 3/5 there numerator and denominator each diff by 2 so i multiplied 5/6 by 2/2 and got 10/12 and then seen well this fits everything else now

OpenStudy (anonymous):

Oh, I see now. I have another question. What would be the pattern for 2+5/2+3+7/2+...+5/6?

OpenStudy (freckles):

well i see the 5/2 and the 7/2 thingy

OpenStudy (anonymous):

*Sorry, the last one is 19/2

OpenStudy (freckles):

those fractions look the most similar out the thingy

OpenStudy (freckles):

so I'm going to see if i can put everything in that form

OpenStudy (freckles):

4/2,5/2,6/2,7/2,...,19/2

OpenStudy (freckles):

4/2 is 2 6/2 is 3

OpenStudy (anonymous):

Oh, I see. Thank you. That's all my questions :)

OpenStudy (freckles):

so the bottom is always going to be 2 and the top thingy starts at 4 and ends at 19 there are (19-4+1) numbers 16 numbers I believe

OpenStudy (freckles):

anyways no problem

OpenStudy (anonymous):

Oh, how were you able to get the numbers?

OpenStudy (anonymous):

\[\sum_{1}^{n?}n/(n+2)\]

OpenStudy (freckles):

you have 4,5,6,7,8,9,10,11,12,13,14,15,16,17,18,19 you can count them or do 19-4+1 like pretend you wanted to know how many numbers (inclusive) is between 5 and 12 there are 12-5+1 if you wanted to know how many numbers are between (inclusive) n and k (n>k) you do n-k+1

OpenStudy (anonymous):

Okay, I get it now. Thank you!

OpenStudy (freckles):

well good luck on your setups let me know if you want me to check something

OpenStudy (anonymous):

the sigma signifies the summation. the lower value below the sigma sign signifies from where the value starts, and the upper limit signifies where it ends (I mistakenly wrote it as n, in place of 'i'). The equation n/(n+2) can be seen from the values given, 1/3, 2/4,3/5...

OpenStudy (anonymous):

Oh, btw going back to the 1st equation. I tried to plug it in and I got 5/7 for the last part.

OpenStudy (freckles):

for the first question you have 10 numbers

OpenStudy (anonymous):

Oh, no. Like the last number for the equation when I tried i/ i +2

OpenStudy (freckles):

\[\sum_{i=1}^{10}\frac{i}{i+2}\] the last number should give you 10/12

OpenStudy (freckles):

which is 5/6

OpenStudy (anonymous):

Oh, nvm. Sorry, my mistake.

OpenStudy (freckles):

you only went to n=5 didn't you? :p

OpenStudy (anonymous):

Yeah >_<

OpenStudy (freckles):

knowing the number of numbers is important to figure out what number we should put on top of that sigma thingy

OpenStudy (anonymous):

Okay. Thank you! You're a really great help :)

OpenStudy (freckles):

ahhh thanks